I had 10 problems on this.
i was able to complete all except 2.
e^2x - 3e^x + 2 = 0
i believe i am supposed to move the 2 to the other side.
but im not quite sure how to solve this problem
id greatly appreciate ur help

Answer by nerdybill(5404)

Apply log rules:
.
Assuming the 2x is both part of the exponent then,
e^2x - 3e^x + 2 = 0
e^2x - e^x^3 + 2 = 0
e^2x - e^x^3 = -2
e^(2x/x^3) = -2
2x/x^3 = e^(-2)
2/x^2 = e^(-2)
2 = e^(-2)x^2
2/e^(-2) = x^2
14.778 = x^2
3.844 = x